solve for x: log(x + 2) + log(x-1)=1
Answer by stanbon(75887)

solve for x: log(x + 2) + log(x-1)=1
-----
log[(x+2)(x-1)] = 1
---
(x+2)(x-1) = 10^1
x^2 + x - 2 = 10
x^2 + x -12 = 0
(x+4)(x-3) = 0
x = -4 or x = 3
------
Only x=3 is a true solution as log(x+2) and log(x-1) would
be meaningless if x = -4.
================================
Cheers,
Stan H.